OpenStudy (howard-wolowitz):

Math question:

OpenStudy (howard-wolowitz):

Questions:

OpenStudy (howard-wolowitz):

Chart:

OpenStudy (lochana):

for first question. you must sum up all numbers in your table

OpenStudy (michele_laino):

you have to fill the "Totals" row, first

OpenStudy (lochana):

second one: republicans under 30 is 196. divide it by the answer of first question. and then multiply by 100

OpenStudy (howard-wolowitz):

ok so : 1st bottom row 459 2nd bottom row 420 3rd bottom row 231 then going down 1st side row 423 2nd side row 341 3rd side row 346 finally the yellow box 1110

OpenStudy (howard-wolowitz):

ok so i got the values but i dont get what i put as the first answer

OpenStudy (michele_laino):

total republican are: \(196+120+143=...\), so the requested percentage, is: \[\frac{{30}}{{196 + 120 + 143}} \cdot 100 = ...\% \]

OpenStudy (howard-wolowitz):

ok i got 6.53%

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

similarly for questions #2 and #3

OpenStudy (michele_laino):

oops.. for questions #3 and #4

OpenStudy (howard-wolowitz):

ok can you lay out the second one for me im sorry i jsut dont wont to do it wrong again

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

question #3: total democrats are: \(138+161+121=...?\), so the requested percentage is: \[\frac{{138}}{{138 + 161 + 121}} \cdot 100 = ...\% \]

OpenStudy (michele_laino):

oops.. I meant total numbers of democrats is...

OpenStudy (michele_laino):

number*

OpenStudy (howard-wolowitz):

what about question 2

OpenStudy (michele_laino):

question #2 is 6.53%

OpenStudy (howard-wolowitz):

so the first one we did was the republican one?

OpenStudy (howard-wolowitz):

then I mean question one

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

the first question is different then the second one

OpenStudy (michele_laino):

question #1 is 1110

OpenStudy (howard-wolowitz):

How many people does the data describe. which is 1110?

OpenStudy (howard-wolowitz):

oh ok lol

OpenStudy (michele_laino):

question #3, is: \[\frac{{138}}{{138 + 161 + 121}} \cdot 100 = ...\% \]

OpenStudy (howard-wolowitz):

32.85%

OpenStudy (michele_laino):

correct!

OpenStudy (howard-wolowitz):

sweet then that makes the last percent around 22% right?

OpenStudy (michele_laino):

I got another result

OpenStudy (howard-wolowitz):

can you show me the lay out then. sorry im an idiot

OpenStudy (michele_laino):

here is my computation: \[\frac{{89}}{{89 + 60 + 82}} \cdot 100 = ...\% \]

OpenStudy (howard-wolowitz):

38.528 or 52.53

OpenStudy (michele_laino):

I got \(38.53 \%\)

OpenStudy (howard-wolowitz):

let me re-work it plz one more time srry

OpenStudy (michele_laino):

ok!

OpenStudy (howard-wolowitz):

i got 38.53% as well

OpenStudy (howard-wolowitz):

may i show you one more?

OpenStudy (michele_laino):

Sorry I have made an error, here are the right percentages: Republicans: 42.70% Democrats: 32.85% Independent: 38.53%

OpenStudy (michele_laino):

hint: what is: \(42.7+32.85+38.53=...?\) namely the sum of the above percentages

OpenStudy (howard-wolowitz):

114.08

OpenStudy (michele_laino):

is it possible to get a percentage greater than 100 % ?

OpenStudy (howard-wolowitz):

no but thats what that add up to?

OpenStudy (howard-wolowitz):

am i not right?

OpenStudy (michele_laino):

please wait...

OpenStudy (howard-wolowitz):

ok

OpenStudy (michele_laino):

we have 423 people under 30, and so we get the subsequent percentages: percentage of republicans people under 30 is: \((196/423)⋅100=...\%\) percentage of democrats people under 30 is: \((138/423)⋅100=...\%\) percentage of independent people under 30 is: \((89/423)⋅100=...\% \)

OpenStudy (michele_laino):

I think that it is what we can say about political party for people under 30

OpenStudy (howard-wolowitz):

were gonna be getting decimals here

OpenStudy (howard-wolowitz):

the first one is 46.33

OpenStudy (michele_laino):

please keep in mind that they are percentages

OpenStudy (michele_laino):

correct!

OpenStudy (howard-wolowitz):

so basically 42.70% is wrong for the answer? of my second question and the correct answer is 46.33%

OpenStudy (howard-wolowitz):

doing these how are we gonna gt the answer to this

OpenStudy (michele_laino):

please note that the first percentages refer to whole people, whereas the second ones refer to people under 30

OpenStudy (michele_laino):

oops.. refer to all people not whole people

OpenStudy (howard-wolowitz):

ok i see .. again im sorry but i am confusing myself here

OpenStudy (michele_laino):

furthermore, we can observe this: since the first percentage about democrats is 32.85, whereas the second percentage is 32.6, then we can say that democrats party is composed by almost people under 30

OpenStudy (michele_laino):

namely the major part of democrats people is under 30

OpenStudy (howard-wolowitz):

so your saying that it would be major part of democarts is under 30

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

do you think that my teacher would understand that for an answer?

OpenStudy (michele_laino):

I think so, since it is logically correct!

OpenStudy (howard-wolowitz):

ok sweet.

OpenStudy (michele_laino):

ok! :)